OpenStudy (anonymous):

Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be _____. A. 14, 28 B. 15, 27 C. 9, 33 D. -13, 55 @phi

OpenStudy (anonymous):

@Johnbc

OpenStudy (anonymous):

@Johnbc

OpenStudy (johnweldon1993):

Let the first integer be 'x' And the second be 'y' " one is three less than 4 times the other" Let 'x' be 'x' And so 'y' will be represented as "4x - 3" "sum of the two integers is 42" So \[\large x + y = 42\] we know y = 4x -3 so \[\large x + 4x - 3 = 42\] Add 3 to both sides \[\large x + 4x = 45\] Add the 'x's \[\large 5x = 45\] what does 'x' = ?

OpenStudy (anonymous):

x= 9?

OpenStudy (anonymous):

Correct

OpenStudy (johnweldon1993):

Right, And if x = 9 Then y (since it is represented as 4x - 3) will equal \[\large y = 4(9) - 3\] \[\large y = ?\]

OpenStudy (anonymous):

so the answer is c? 9,33?

OpenStudy (anonymous):

You can use your x value to solve for your Y value or look at your answer choices and see if you can already identify a match.

OpenStudy (anonymous):

Correct

OpenStudy (johnweldon1993):

Correct!

OpenStudy (anonymous):

thank you both!!